# The Open Journal of Astrophysics

### v2

@@ Acknowledgements
We thank Andrew Hearin and Katrin Heitmann for helpful conversations. S.P. thanks Becky Nevin and Aleksandra Ciprajanovic for useful conversations regarding uncertainty quantification. S.P. acknowledges support from the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, https://iaifi.org). Y.Y. acknowledges support from the Khoury Apprenticeship program. J.B. and N.V.A. are supported in this work by NSF award AST-2206563, the US DOE under grant DE-SC0024787, and the Roman Research and Support Participation program under NASA grant 80NSSC24K0088. R.W. is supported by NSF award DMS-2134178. Data generation was conducted on the Discovery cluster, supported by Northeastern University’s Research Computing team. The machine learning computations were run on the FASRC cluster supported by the FAS Division of Science Research Computing Group at Harvard University.
@@ Appendix A HOD Model Construction
Constructing an HOD model with alignment requires us to choose an occupation model component to populate dark matter halos, a phase space model component to place these galaxies within their halo, and an alignment model component to assign an orientation to the galaxies. Since we distinguish between central and satellite galaxies, we choose components for each of these steps for each of the two populations. We choose the Zheng07Cens and Zheng07Sats occupation model components for central and satellite galaxies, respectively. These components are available in halotools and implement Equations 2, 3 and 5 from Zheng et al. (2007). As discussed earlier, the five HOD parameters chosen for these models come from Table 1 of the same paper.
For simplicity, we use TrivialPhaseSpace and SubhaloPhaseSpace as phase space model components for the central and satellite galaxies respectively, which places central galaxies at the location of their parent halo and satellite galaxies at the location of any subhalos (smaller dark matter halos that reside within the larger parent halo). We use CentralAlignment and RadialSatelliteAlignment for the alignment model components. The CentralAlignment component aligns the central galaxy with respect to its parent halo, and RadialSatelliteAlignment aligns the satellite galaxies with respect to the radial vector between the central galaxy and itself. The parameters $\mu_{\rm cen}$ and $\mu_{\rm sat}$, the central and satellite alignment strengths, determine the shape of the Dimroth-Watson distribution from which the misalignment angles are drawn (Van Alfen et al., 2024).
@@ Appendix B Correlation rescaling
As a result of taking the $\log$ of $\xi$ for training, the rescaled aleatoric and epistemic uncertainties are only well-defined in log space:
$\sigma_{\log\xi}^{\text{aleo}}=\sigma_{\log\bar{\xi}}\cdot\widehat{\sigma_{\log\xi}^{\text{aleo}}},\qquad\sigma_{\log\xi}^{\text{epi}}=\sigma_{\log\bar{\xi}}\cdot\widehat{\sigma_{\log\xi}^{\text{epi}}}.$
@@
$\omega=\omega^{\prime}\cdot\sigma_{\bar{\omega}}+\mu_{\bar{\omega}},\qquad\sigma_{\omega}^{\text{aleo}}=\sigma_{\bar{\omega}}\cdot\widehat{\sigma_{\omega}^{\text{aleo}}},\qquad\sigma_{\omega}^{\text{epi}}=\sigma_{\bar{\omega}}\cdot\widehat{\sigma_{\omega}^{\text{epi}}}.$
The last correlation function $\eta$ is rescaled similarly to $\omega$.
Appendix C Training and Hamiltonian Monte Carlo Details
Training. We train the IAEmu for a maximum of 500 epochs with a 100-epoch warm-up period and early stopping. We also employ gradient clipping for numerical stability, as the training of MVE networks can suffer from instability. The use of residual connections and a shallower embedding network than the decoder is to stabilize convergence during training. We employ various techniques to further aid the convergence of the model. Following the recommendations in (Sluijterman et al., 2023), we initialize all variance output-neurons to have a bias of zero which results in a constant variance prediction across all bins at initialization, ensuring that no bins are biased towards large variances. We additionally implement a warm-up period during training with $\beta=1.0$ for all correlations to maximize regression on the means before transitioning to a value of $\beta_{\xi}=0.9$ and $\beta_{\omega}=\beta_{\eta}=0.5$ for the remainder of training. The value of $\beta_{\xi}$ was chosen as $\xi(r)$ correlations exhibit a very high signal-to-noise ratio, so the aleatoric uncertainties on these correlations are generally not significant or of interest.
We use the AdamW optimizer (Loshchilov & Hutter, 2019) with a training batch size of 128 and a step learning rate scheduler (10% decay at 167-epoch intervals with a starting $\text{lr}=0.01$). Additional L2-regularization via a weight decay factor of $10^{-4}$ is used in the optimizer. All training was done on two NVIDIA A100-80GB GPUs. During training, IAEmu is validated every 5 epochs with an early stopping patience of 100 epochs based on the validation criteria. The validation criteria for saving the model is a linear combination of MSE and Gaussian-NLL losses computed for each correlation $\xi$, $\omega$, and $\eta$.
@@ Appendix C Training and Hamiltonian Monte Carlo Details
Equation 11 is a form of Bayes’ Theorem where $p(\mathcal{D}|x,\theta)$ is the likelihood function and $p(x)$ is the prior distribution on $x$. HMC achieves this by forward modeling the dynamics of a governing Hamiltonian $H$:
@@
where T is the kinetic energy with mass matrix $M$ and momentum $p$, which controls the exploration in parameter space, and $-\ln p(x|\mathcal{D},\theta)$ takes the role of the potential energy $U$. The time-evolution of $x$ and $p$ is accordingly governed by Hamilton’s equations. HMC thus arrives at the posterior distribution over the inputs by sequentially evolving the dynamical variables according to Hamiltonian dynamics; this of course corresponds to minimizing the potential energy, which maximizes the log probability. Hamilton’s equations require gradients with respect to the Hamiltonian $H$, specifically $-\nabla_{x}\ln p(x|\mathcal{D},\theta)$. Decomposing this with chain rule,
@@ Appendix D Example IAEMU Predictions
Randomly chosen IAEmu test-set predictions with the halotools-IA ground truth shown in blue, and IAEmu with $1\sigma$ epistemic uncertainties given in red.
Figure 9: 25 random IAEmu test set predictions for $\xi$. $\log(\xi)$ is plotted due to the large range in correlation amplitudes. Error bars on ground truth values are computed across 10 realizations of halotools-IA. $1\sigma$ epistemic uncertainty is shown in the red shaded region.
Figure 10: 25 random IAEmu test set predictions for $\omega$. Error bars on ground truth values are computed across 10 realizations of halotools-IA. $1\sigma$ epistemic uncertainty is shown in the red shaded region.
Figure 11: 25 random IAEmu test set predictions for $\eta$. Error bars on ground truth values are computed across 10 realizations of halotools-IA. $1\sigma$ epistemic uncertainty is shown in the red shaded region.

### v4

@@ Acknowledgements
We thank Andrew Hearin and Katrin Heitmann for helpful conversations. S.P. thanks Becky Nevin and Aleksandra Ćiprijanović for useful conversations regarding uncertainty quantification. S.P. acknowledges support from the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, https://iaifi.org). Y.Y. acknowledges support from the Khoury Apprenticeship program. J.B., N.V.A. and Y.Y. are supported in this work by NSF award AST-2206563, the US DOE under grant DE-SC0024787, and the Roman Research and Support Participation program under NASA grant 80NSSC24K0088. R.W. is supported by NSF award DMS-2134178. Data generation was conducted on the Discovery cluster, supported by Northeastern University’s Research Computing team. The machine learning computations were run on the FASRC cluster supported by the FAS Division of Science Research Computing Group at Harvard University.
@@ Appendix A HOD Model Construction
Constructing an HOD model with alignment requires us to choose an occupation model component to populate dark matter halos, a phase space model component to place these galaxies within their halo, and an alignment model component to assign an orientation to the galaxies. Since we distinguish between central and satellite galaxies, we choose components for each of these steps for each of the two populations. We choose the Zheng07Cens and Zheng07Sats occupation model components for central and satellite galaxies, respectively. These components are available in halotools and implement Equations 2, 3 and 5 from Zheng_2007. As discussed earlier, the five HOD parameters chosen for these models come from Table 1 of the same paper.
For simplicity, we use TrivialPhaseSpace and SubhaloPhaseSpace as phase space model components for the central and satellite galaxies respectively, which places central galaxies at the location of their parent halo and satellite galaxies at the location of any subhalos (smaller dark matter halos that reside within the larger parent halo). We use CentralAlignment and RadialSatelliteAlignment for the alignment model components. The CentralAlignment component aligns the central galaxy with respect to its parent halo, and RadialSatelliteAlignment aligns the satellite galaxies with respect to the radial vector between the central galaxy and itself. The parameters $\mu_{\rm cen}$ and $\mu_{\rm sat}$, the central and satellite alignment strengths, determine the shape of the Dimroth-Watson distribution from which the misalignment angles are drawn (vanalfen_2023).
@@ Appendix B Correlation rescaling
As a result of taking the $\log$ of $\xi$ for training, the rescaled aleatoric and epistemic uncertainties are only symmetric in log space. The corresponding transformation of uncertainties from log to linear space follows from the log normal moments:
$\sigma_{\xi}^{\text{aleo}}$ $=\sqrt{\left(e^{(\sigma_{\log\xi}^{\text{aleo}})^{2}}-1\right)e^{2(\log\hat{\xi}\,\sigma_{\log\bar{\xi}}+\mu_{\log\bar{\xi}})+(\sigma_{\log\xi}^{\text{aleo}})^{2}}},$
$\sigma_{\xi}^{\text{epi}}$ $=\exp\!\big(\log\hat{\xi}\,\sigma_{\log\bar{\xi}}+\mu_{\log\bar{\xi}}\big)\cdot\sigma_{\log\bar{\xi}}\cdot\widehat{\sigma_{\log\xi}^{\text{epi}}}$.
Here the aleatoric term follows from the exact log normal variance, while the epistemic term is obtained by linear error propagation through the exponential mapping.
Figure 9: Aleatoric vs. epistemic uncertainty comparison for $\xi$ with correlation uncertainty bias (left) and correlation amplitude bias (right). For test-set predictions, we analyze the total spread of aleatoric uncertainties of the data predicted by IAEmu and epistemic uncertainties due to the stochasticity of IAEmu. In the left plot, the coloring corresponds to the log-residual between IAEmu predicted aleatoric uncertainties and (true) aleatoric uncertainties from halotools-IA produced from the 10 realizations used in producing the dataset. In the right plot, the coloring corresponds to the log-residual between IAEmu predicted correlation amplitudes and (mean) ground truth amplitudes from halotools-IA produced from 10 realizations used in producing the dataset. It is seen that there is no clear correlations between log-residuals in the correlation amplitudes and IAEmu aleatoric and epistemic uncertainties. It is also seen that aleatoric uncertainties for $\xi$ are biased high when epistemic uncertainties are largest.
We perform an analogous uncertainty analysis for $\xi$, following the methodology applied to $\omega$ and $\eta$ in Section 4.2. The results are presented in Figure 9. This analysis was omitted from the main discussion because $\xi$ exhibits a significantly higher signal correlation than $\omega$ and $\eta$, and the IAEmu training procedure was not optimized to prioritize aleatoric uncertainty estimation for this quantity (see Section 3.2). Nonetheless, the findings parallel those obtained for $\omega$ and $\eta$: the IAEmu architecture appears sufficiently expressive for this task, as the scatter consistently lies below the 1:1 line. We observe a median bias of $0.38$ dex in the aleatoric uncertainties, which exceeds the biases observed for $\omega$ and $\eta$. This discrepancy likely stems from the choice of $\beta_{\xi}$, which was not tuned to facilitate accurate uncertainty quantification for $\xi$. Consequently, IAEmu tends to overestimate the aleatoric uncertainty when the epistemic uncertainty is large. In contrast, a negligible median bias of $0.01$ dex is found for the correlation amplitudes.
@@
$\omega=\omega^{\prime}\cdot\sigma_{\bar{\omega}}+\mu_{\bar{\omega}},\quad\sigma_{\omega}^{\text{aleo}}=\sigma_{\bar{\omega}}\cdot\widehat{\sigma_{\omega}^{\text{aleo}}},\quad\sigma_{\omega}^{\text{epi}}=\sigma_{\bar{\omega}}\cdot\widehat{\sigma_{\omega}^{\text{epi}}}$
$\eta=\eta^{\prime}\cdot\sigma_{\bar{\eta}}+\mu_{\bar{\eta}},\quad\sigma_{\eta}^{\text{aleo}}=\sigma_{\bar{\eta}}\cdot\widehat{\sigma_{\eta}^{\text{aleo}}},\quad\sigma_{\eta}^{\text{epi}}=\sigma_{\bar{\eta}}\cdot\widehat{\sigma_{\eta}^{\text{epi}}}$.
Appendix C Training and Hamiltonian Monte Carlo
Training. We train the IAEmu for a maximum of 500 epochs with a 100-epoch warm-up period and early stopping. We also employ gradient clipping for numerical stability, as the training of MVE networks can suffer from instability. The use of residual connections and a shallower embedding network than the decoder is to stabilize convergence during training. We employ various techniques to further aid the convergence of the model. Following the recommendations in (sluijterman2023optimaltrainingmeanvariance), we initialize all variance output-neurons to have a bias of zero which results in a constant variance prediction across all bins at initialization, ensuring that no bins are biased towards large variances.
The choice of the $\beta$ parameter in the $\beta$-NLL loss (equation 7) dictates the strength in which the loss interpolates between standard MSE and Gaussian-NLL loss. The optimal value of $\beta$ will also not be the same for each correlation that is predicted. We implement a warm-up period during training with $\beta=1.0$ for all correlations to maximize regression on the means before transitioning to a value of $\beta_{\xi}=0.9$ and $\beta_{\omega}=\beta_{\eta}=0.5$ for the remainder of training. The value of $\beta_{\xi}$ was chosen as $\xi(r)$ correlations exhibit a very high signal-to-noise ratio, so the aleatoric uncertainties on these correlations are generally not significant or of interest.
We use the AdamW optimizer (loshchilov2019decoupled) with a training batch size of 128 and a step learning rate scheduler (10% decay at 167-epoch intervals with a starting learning rate of 0.01). Additional L2-regularization via a weight decay factor of $10^{-4}$ is used in the optimizer. All training was done on two NVIDIA A100-80GB GPUs. During training, IAEmu is validated every 5 epochs with an early stopping patience of 100 epochs based on the validation criteria. The validation criteria for saving the model is a linear combination of MSE and Gaussian-NLL losses computed for each correlation $\xi$, $\omega$, and $\eta$.
@@ Appendix C Training and Hamiltonian Monte Carlo
Equation 13 is a form of Bayes’ Theorem, where $p(\mathcal{D}|x,\theta)$ is the likelihood function and $p(x)$ is the prior distribution on $x$. HMC achieves this by forward modeling the dynamics of a governing Hamiltonian $H$:
@@
where $T$ is the kinetic energy with mass matrix $M$ and momentum $\mathbf{p}$, which controls the exploration in parameter space, and $\ln p(x|\mathcal{D},\theta)$ takes the role of the potential energy $U$. The time-evolution of $x$ and $p$ is accordingly governed by Hamilton’s equations:
$\frac{dx}{dt}=\frac{\partial H}{\partial\mathbf{p}},\qquad\frac{d\mathbf{p}}{dt}=-\frac{\partial H}{\partial x}.$ (15)
HMC thus arrives at the posterior distribution over the inputs by sequentially evolving the dynamical variables according to Hamiltonian dynamics; this of course corresponds to minimizing the potential energy, which maximizes the log probability. As seen in equation 15, Hamilton’s equations require gradients with respect to $H$, specifically $\nabla_{x}\ln p(x|\mathcal{D},\theta)$. Decomposing this with chain rule,
@@ Appendix D Example IAEmu Predictions
Randomly chosen IAEmu test-set predictions comparing halotools-IA and IAEmu are shown in Figures 10, 11, and 12. In the figures, the halotools-IA ground truth is shown in blue, and IAEmu with $1\sigma$ epistemic uncertainties are given in red.
Figure 10: 25 random IAEmu test set predictions (red) compared with halotools-IA ground truth (blue) for $\xi$. $\log(\xi)$ is plotted due to the large range in correlation amplitudes. Error bars on ground truth values are computed across 10 realizations of halotools-IA. $1\sigma$ epistemic uncertainty for IAEmu is shown in the red shaded region. The general agreement between IAEmu and halotools-IA for $\xi$ is observed.
Figure 11: 25 random IAEmu test set predictions (red) compared with halotools-IA ground truth (blue) for $\omega$. Error bars on ground truth values are computed across 10 realizations of halotools-IA. $1\sigma$ epistemic uncertainty for IAEmu is shown in the red shaded region. The general agreement between IAEmu and halotools-IA for $\omega$ is observed, as well as the stochasticity, zero-crossings, and small amplitudes of the correlation. As mentioned in Section 4.1, these features make quantifying IAEmu performance in terms of fractional error difficult.
Figure 12: 25 random IAEmu test set predictions (red) compared with halotools-IA ground truth (blue) for $\eta$. Error bars on ground truth values are computed across 10 realizations of halotools-IA. $1\sigma$ epistemic uncertainty for IAEmu is shown in the red shaded region. The general agreement between IAEmu and halotools-IA for $\eta$ is observed, as well as the stochasticity, zero-crossings, and small amplitudes of the correlation. As mentioned in Section 4.1, these features make quantifying IAEmu performance in terms of fractional error difficult.